# Multi-Level Feature Aggregation and Recursive Alignment Network for Real-Time Semantic Segmentation

### v1

@@ IV-B3 Ablation Study for RAM
IV-B4 Ablation Study for ASFM
We design ASFM to fuse multi-scale scores in an adaptive manner. Three fusion methods are compared in Table. IV: Average, Maximum (Max), and ASFM. When using Average or Max operations, the score maps at each scale are either equally important or sparsely selected, making them a suboptimal choice. ASFM improves the segmentation accuracy by more than 1.0% over the other two methods, demonstrating the effectiveness of this adaptive fusion strategy. Compared with the promotion of accuracy, the increase in computational complexity is acceptable, as the inference speed is only reduced by around 5 FPS and remains above the real-time speed.
[…]
TABLE IX: Comparison with real-time methods on the Cityscapes dataset. In addition to using the whole picture as input to compute mIoU, we also provide single-scale inference accuracy on the crop size 1024 $\times$ 1024. ’MFARANet-Pruned’ indicates the pruning method specific to MFARANet, which is shown in Fig. 8 and discussed in Section. IV-B8.
Year Name mIoU(%) Params(M) Resolution
@@ IV-B4 Ablation Study for ASFM
2016 ENet [57] - 58.3 0.4 360 $\times$ 640
2018 ESPNet [58] - 60.3 0.4 512 $\times$ 1024
2018 BiSeNetV1-1 [48] 69.0 68.4 5.8 768 $\times$ 1536
2018 ContextNet [59] - 66.1 0.9 1024 $\times$ 2048
2019 Fast-SCNN [60] 68.6 68.0 1.1 1024 $\times$ 2048
2019 DFANetB [20] - 67.1 4.8 1024 $\times$ 1024
2019 EDANet [61] - 67.3 0.7 512 $\times$ 1024
2020 CGNet_M3N21 [62] - 64.8 0.5 360 $\times$ 640
2020 CIFReNet [63] 72.9 70.9 1.9 512 $\times$ 1024
2020 NDNet [34] - 65.7 1.1 1024 $\times$ 2048
2020 MiniNet-V2 [64] - 70.5 0.5 512 $\times$ 1024
2020 AGLNet [65] - 70.1 1.1 512 $\times$ 1024
2022 SGCPNet2 [37] - 69.5 0.6 768 $\times$ 1536
@@
2017 SegNet [40] - 57.0 29.5 360 $\times$ 640
2017 ERFNet [8] 70.0 68.0 20.0 512 $\times$ 1024
2018 ICNet [9] - 69.5 26.5 1024 $\times$ 2048
2018 BiSeNetV1-2 [48] 74.8 74.7 49.0 768 $\times$ 1536
2019 SwiftNet-18 [13] 75.5 75.4 11.8 1024 $\times$ 2048
2019 SwiftNet-18pyr [13] 74.4 75.1 12.9 1024 $\times$ 2048
2019 ShelfNet [35] - 74.8 23.5 512 $\times$ 512
2020 BiSeNetV2-L [6] 75.8 75.3 - 512 $\times$ 1024
2020 SFNet(DF1) [36] - 74.5 9.0 1024 $\times$ 2048
2021 STDC2-Seg75 [66] 77.0 76.8 12.5 768 $\times$ 1536
2021 RGPNet [23] 74.1 - 17.8 1024 $\times$ 2048
2021 FaPN2 [42] 75.6 75.0 12.6 768 $\times$ 1536
2022 DMA-Net [21] - 75.6 14.6 768 $\times$ 1536
@@
IV-B5 Importance of each Module
In Table. V, we conduct three experiments to reveal the importance of feature aggregation, spatial alignment and adaptive score-level fusion. In the first row, we directly input the features from each level of the encoder into subsequent modules. This results in a significant reduction in accuracy of 2.6%, suggesting that aggregating multi-level features at each scale can benefit spatial alignment and inference. In the second experiment, we utilize bilinear interpolation to replace RAM for upsampling low-resolution features, which decreases the accuracy by 1.5%. This result reveals that spatial misalignment is the key factor limiting scores fusion, and our RAM can effectively address this problem. Finally, we use concatenation operation followed by a 3 $\times$ 3 convolutional layer to fuse multi-scale features first, and then use a segmentation head to obtain the prediction. We conduct this experiment to compare the commonly used feature-level fusion with our in-network score-level fusion. Our ASFM reduces the computational complexity by 17.2 GFLOPs while improving the accuracy by 0.6%, which proves the effectiveness and efficiency of the adaptive score-level fusion mechanism.
TABLE X: Comparison with accuracy-oriented methods on the Cityscapes test set. All models are trained using only the finely annotated dataset and the results are reported using multi-scale testing.
 Name Backbone mIoU(%) Params(M)
FCN8s [4] VGG 65.3 50.5
DeepLab-V2 [32] ResNet-101 70.4 44.0
RefineNet [67] ResNet-101 73.6 118.4
SAC [68] ResNet-101 78.1 -
DepthSeg [69] ResNet-101 78.2 -
PSPNet [11] ResNet-101 78.4 65.7
ResNet38 [70] - 78.4 -
BiSeNet-V1 [48] ResNet-101 78.9 51.0
DFN [71] ResNet-101 79.3 90.7
PSANet [72] ResNet-101 80.1 85.6
DANet [73] ResNet-101 81.5 66.6
HRNet-V2 [5] HRNetV2-W48 81.6 65.9
MFARANet ResNet-18 78.7 13.0
TABLE XI: Model complexity comparison of representative models in Table. IX and Table. X. FLOPs and FPS are computed on a single GTX 3090 GPU with an input size of 1024 $\times$ 1024.
 Name FLOPs(G)↓ Speed(FPS)↑ Params(M)↓
Speed-Oriented:
ENet [57] 10.0 53.4 0.4
ESPNet [58] 10.1 80.3 0.4
ContextNet [59] 21.5 66.4 0.9
Fast-SCNN [60] 3.1 166.2 1.1
DFANetB [20] 7.1 40.1 4.8
EDANet [61] 16.5 94.8 0.7
CGNet_M3N21 [62] 12.6 70.1 0.5
CIFReNet [63] 30.2 49.3 1.9
MiniNet-V2 [64] 17.2 63.9 0.5
AGLNet [65] 30.0 42.6 1.1
SGCPNet2 [37] 1.5 66.9 0.6
Good Trade-Off:
ERFNet [8] 55.5 44.2 20.0
ICNet [9] 137.2 29.9 26.5
SwiftNet-18 [13] 48.2 117.6 11.8
SwiftNet-18pyr [13] 59.7 73.0 12.9
ShelfNet [35] 46.4 114.1 23.5
SFNet(DF1) [36] 15.2 140.1 9.0
STDC2-Seg75 [66] 75.4 70.2 12.5
DMA-Net [21] 44.2 93.2 14.6
Accuracy-Oriented:
DeepLab-V2 [32] 1236.8 4.1 44.0
RefineNet [67] 979.1 9.8 118.4
PSPNet [11] 696.2 8.5 65.7
BiSeNet-V1 [48] 193.4 25.8 51.0
DFN [71] 1006.3 8.2 90.7
HRNet-V2 [5] 348.1 14.1 65.9
MFARANet-Pruned 81.3 64.4 12.2
MFARANet 101.6 52.2 13.0
IV-B6 Ablation Study for Booster Training Strategy
Figure 9: Inference speed and Params vs. mIoU accuracy on the Cityscapes test set. The size of the points indicates the model size. Purple points are representative real-time methods in Table. IX. The red and green points represent our MFARANet using 1024 $\times$ 1024 or the whole image (1024 $\times$ 2048) input for inference, respectively. All FPS are measured on a single GTX 3090 GPU with the image resolution on which the inference is performed to calculate the accuracy.
Table. VI gives the performance comparison of MSL and MJS with and without OHEM, respectively. Compared to MSL, our MJS brings a 0.9% reduction in accuracy. This may be due to the fact that directly adding boundary supervision at each scale will lead to a more serious class imbalance problem, since boundary pixels only occupy a small part of the whole image. Boundary pixels are difficult training examples, so we employ OHEM to training on hard pixels selected by current loss to alleviate the class imbalance problem. After combing with OHEM, the accuracy of MJS + OHEM exceeds MSL + OHEM. The accuracy increase of MJS is 2.1%, while MSL only improves by 0.1%. This proves that our MJS is highly complementary to OHEM, and using them together maximizes the accuracy of our model. By adopting the booster training strategy, our model finally achieves a segmentation accuracy of 78.2% with real-time speed on the validation set of Cityscapes.
IV-B7 Ablation for Multi-scale Predictions
@@
Furthermore, we conduct several ablation experiments in Table. VII on how multi-scale maps contribute to the final results. When removing the score map at a certain scale, we additionally add a segmentation head and a boundary head to calculate auxiliary loss, ensuring that all ablations can use the same loss MJS+OHEM introduced in Section. III-E. This also helps maintain the structural integrity of the MFAM and RAM. The results show that training on a single-scale map reduces the accuracy by 0.96%-1.65% with a promotion of only 4.7 FPS, while discarding one of the four score maps results in a decrease of 0.49%-1.04% with almost the same model complexity. Therefore, we can conclude that training on full-scale score maps is necessary as it achieves a better trade-off than the above ablations.
IV-B8 Ablation for Scale Selection
@@
IV-C Comparison With State-of-the-Arts on the Cityscapes
IV-C1 Accuracy Comparison
We compare our method with other classical or newly proposed real-time semantic segmentation methods in Table. IX. We roughly split them into two categories: the first one is speed-oriented, which has extremely low model complexity (FLOPs or Params); the other achieves a good trade-off and can reach relatively high accuracy with moderate model complexity. Compared to these speed-oriented method, our network improves segmentation accuracy by at least 6.4% on the test dataset while maintaining real-time speed on a relatively high-resolution image of 1024 $\times$ 2048. In particular, we outperform the well-known ENet by a large margin of 19.0%. As for the balance-oriented methods, they improve the accuracy as much as possible while keeping a real-time inference speed of over 24 FPS. SwiftNet and SFNet achieve close accuracy to our method, but both of them employ modified versions of Pyramid Pooling Module (PPM) [11] on top of their backbone to obtain multi-scale features, significantly improving their performance. Overall, without using additional modules (e.g., Atrous Spatial Pyramid Pooling [32], PPM) or their modified versions, our method with moderate model size achieves the highest segmentation accuracy among all real-time methods.
TABLE XII: Performance on CamVid. We use the whole picture as input to compute mIoU.
@@ IV-C1 Accuracy Comparison
We also compare our method with accuracy-oriented methods in Table. X, most of which use a very deep classification network (e.g., ResNet-101) or a well-designed large-scale network (e.g., HRNet) as the backbone. Our method shows a very competitive result with these state-of-the-art methods with much smaller model size. Compared to the most accurate method HRNetV2, our network reduces the accuracy by less than 3.0% at 19.7% of the model size, 29.2% of the FLOPs and 3.7$\times$ FPS (Table. XI). Furthermore, by using a shallow backbone, our network reaches a higher segmentation accuracy with a much smaller model size than most methods, including the well-known DeepLab, RefineNet, PSPNet, etc. This shows that our network has a light-weight and effective decoder, and further demonstrates that using the inherent multi-scale property of the network to preform in-network multi-scale inference is one way to achieve efficient semantic segmentation.
IV-C2 Model Complexity Comparison
Table. XI is the model complexity comparisons between real-time and accuracy-oriented methods, from which we have the following observations: 1) Our MFARANet achieves inference speeds of 52.2 and 64.4 FPS (with pruning) on a relatively large input size of 1024 $\times$ 1024 at medium model sizes. These speeds are much higher than the real-time threshold of 24 FPS, indicating that our model has the potential to achieve real-time speeds on GPUs with lower computing power. 2) We achieve better or competitive performance with much smaller model complexity compared to accuracy-oriented methods. For example, our model outperforms PSPNet by 0.3% mIoU (Table. X) with 14.6% FLOPs, 19.8% Params and 6.2$\times$ FPS of it. 3) The proposed MFARANet has the second largest FLOPs among real-time methods, while the speed of the pruned version is higher than 7 algorithms and close to more than half of the models (difference is less than 10 FPS). This is because some methods adopt inefficient implementations such as convolutional factorization to reduce FLOPs, which increases their actual processing time on the GPU platform.
Fig. 9 reflects the trade-off between FPS and accuracy, from which we can draw the following conclusions: 1) Our MFARANet and its pruned versions using different input size reach the highest accuracy among all real-time methods. 2) Only STDC2-Seg75 could reach competitive accuracy with our model (only 0.3% lower than us) on the test set at a similar inference speed, while lower than us by 0.9% on the valuation set (Table. IX). Additionally, our pruned-model has the same accuracy as STDC2-Seg75 on the test set, and reaches 0.7% higher accuracy on the valuation set with a speed increase of 12.1 FPS. 3) For methods having a faster inference speed, our MFARANet-pruned at 64.4 FPS achieves 1.2%-9.5% accuracy improvement on the test set.
IV-D Evaluation on Other Datasets
IV-D1 Performance on CamVid
TABLE XIII: Performance on PASCAL-Context. All results are reported using multi-scale testing. We also report the performance of our model with different backbones.
@@ IV-D1 Performance on CamVid
CamVid is another dataset of road scene from a car perspective, which has much fewer training and testing samples than Cityscapes. We report the results on it in Table. XII to prove the generality of our approach, from which we can draw similar conclusions as on the Cityscapes dataset. Compared to speed-oriented methods, our method shows a large increase in accuracy. For example, compared to ENet and CGNet, the accuracy of our method improves by 21.6% and 7.3%, respectively. Among methods aiming at good trade-off, we achieve the highest accuracy with a moderate model size.
IV-D2 Performance on PASCAL-Context
@@
PASCAL-Context is a very challenging scene understanding dataset. Quantitative results are shown in Table. XIII, and all comparisons are picked from accuracy-oriented methods. Using ResNet-18 as the encoder, our method can outperform half of the accuracy-oriented methods in the table with very few parameters. By employing ResNet-50, the MFARANet can exceed the performance of all methods except EncNet using ResNet-101, including RefineNet adopting the very deep classification network ResNet-152 as the backbone. In particular, when using the same ResNet-50 as the encoder, our method improves the accuracy by 0.4% and 4% compared to EncNet and ShelfNet50, while having a much smaller model size (25.7M vs. 35.5M and 38.7M). The above results demonstrate the efficiency and effectiveness of our method.
V Conclusion

### v3

@@ 4.2.6 Ablation Study for Booster Training Strategy
Table. 6 gives the performance comparison of MSL and MJS with and without OHEM, respectively. Compared to MSL, our MJS brings a 0.9% reduction in accuracy. This may be due to the fact that directly adding boundary supervision at each scale will lead to a more serious class imbalance problem, since boundary pixels only occupy a small part of the whole image. Boundary pixels are difficult training examples, so we employ OHEM to training on hard pixels selected by current loss to alleviate the class imbalance problem. After combing with OHEM, the accuracy of MJS + OHEM exceeds MSL + OHEM. The accuracy increase of MJS is 2.1%, while MSL only improves by 0.1%. This proves that our MJS is highly complementary to OHEM, and using them together maximizes the accuracy of our model. By adopting the booster training strategy, our model finally achieves a segmentation accuracy of 78.2% with real-time speed on the validation set of Cityscapes.
4.2.7 Ablation for Multi-scale Predictions
Table 9: Comparison with real-time methods on the Cityscapes dataset. In addition to using the whole picture as input to compute mIoU, we also provide single-scale inference accuracy on the crop size 1024 $\times$ 1024. ’MFARANet-Pruned’ indicates the pruning method specific to MFARANet, which is shown in Fig. 9 and discussed in Section. 4.2.8.
 Name mIoU(%) Params(M) Resolution
@@ 4.2.7 Ablation for Multi-scale Predictions
ENet [63] - 58.3 0.4 360 $\times$ 640
ESPNet [64] - 60.3 0.4 512 $\times$ 1024
BiSeNetV1-1 [54] 69.0 68.4 5.8 768 $\times$ 1536
ContextNet [65] - 66.1 0.9 1024 $\times$ 2048
Fast-SCNN [66] 68.6 68.0 1.1 1024 $\times$ 2048
DFANetB [21] - 67.1 4.8 1024 $\times$ 1024
EDANet [67] - 67.3 0.7 512 $\times$ 1024
CGNet_M3N21 [68] - 64.8 0.5 360 $\times$ 640
CIFReNet [69] 72.9 70.9 1.9 512 $\times$ 1024
NDNet [40] - 65.7 1.1 1024 $\times$ 2048
MiniNet-V2 [70] - 70.5 0.5 512 $\times$ 1024
AGLNet [71] - 70.1 1.1 512 $\times$ 1024
SGCPNet2 [43] - 69.5 0.6 768 $\times$ 1536
@@
SegNet [46] - 57.0 29.5 360 $\times$ 640
ERFNet [7] 70.0 68.0 20.0 512 $\times$ 1024
ICNet [8] - 69.5 26.5 1024 $\times$ 2048
BiSeNetV1-2 [54] 74.8 74.7 49.0 768 $\times$ 1536
SwiftNet-18 [12] 75.5 75.4 11.8 1024 $\times$ 2048
SwiftNet-18pyr [12] 74.4 75.1 12.9 1024 $\times$ 2048
ShelfNet [41] - 74.8 23.5 512 $\times$ 512
BiSeNetV2-L [72] 75.8 75.3 - 512 $\times$ 1024
SFNet(DF1) [42] - 74.5 9.0 1024 $\times$ 2048
STDC2-Seg75 [73] 77.0 76.8 12.5 768 $\times$ 1536
RGPNet [24] 74.1 - 17.8 1024 $\times$ 2048
FaPN2 [48] 75.6 75.0 12.6 768 $\times$ 1536
DMA-Net [22] - 75.6 14.6 768 $\times$ 1536
@@
Table 10: Comparison with accuracy-oriented methods on the Cityscapes test set. All models are trained using only the finely annotated dataset and the results are reported using multi-scale testing. FLOPs and FPS are computed on a single GTX 3090 GPU with an input size of 1024 $\times$ 1024.
Name
FLOPs
(G)↓
Speed
(FPS)↑
Params
(M)↓
mIoU
FCN8s [30] 654.9 12.3 50.5 65.3
DeepLab-V2 [38] 1236.8 4.1 44.0 70.4
RefineNet [74] 979.1 9.8 118.4 73.6
SAC [75] - - - 78.1
DepthSeg [76] - - - 78.2
PSPNet [10] 696.2 8.5 65.7 78.4
ResNet38 [77] 2231.1 3.4 - 78.4
BiSeNet-V1 [54] 193.4 25.8 51.0 78.9
DFN [78] 1006.3 8.2 90.7 79.3
PSANet [79] 806.0 - 85.6 80.1
DANet [80] 1052.1 5.4 66.6 81.5
HRNet-V2 [15] 348.1 14.1 65.9 81.6
MFARANet-Pruned 81.3 64.4 12.2 78.5
MFARANet 101.6 52.2 13.0 78.7
Furthermore, we conduct several ablation experiments in Table. 7 on how multi-scale maps contribute to the final results. When removing the score map at a certain scale, we additionally add a segmentation head and a boundary head to calculate auxiliary loss, ensuring that all ablations can use the same loss MJS+OHEM introduced in Section. 3.5. This also helps maintain the structural integrity of the MFAM and RAM. The results show that training on a single-scale map reduces the accuracy by 0.96%-1.65% with a promotion of only 4.7 FPS, while discarding one of the four score maps results in a decrease of 0.49%-1.04% with almost the same model complexity. Therefore, we can conclude that training on full-scale score maps is necessary as it achieves a better trade-off than the above ablations.
4.2.8 Ablation for Scale Selection
@@
4.3 Comparison With State-of-the-Arts on the Cityscapes
4.3.1 Accuracy Comparison
We compare our method with other real-time semantic segmentation methods in Table. 9. We roughly split them into two categories: the first one is speed-oriented, which has extremely low model complexity (FLOPs or Params); the other achieves a good trade-off and can reach relatively high accuracy with moderate model complexity. Compared to these speed-oriented method, our network improves segmentation accuracy by at least 6.4% on the test dataset while maintaining real-time speed on a relatively high-resolution image of 1024 $\times$ 2048. In particular, we outperform the well-known ENet by a large margin of 19.0%. As for the balance-oriented methods, they improve the accuracy as much as possible while keeping a real-time inference speed of over 24 FPS. SwiftNet and SFNet achieve close accuracy to our method, but both of them employ modified versions of Pyramid Pooling Module (PPM) [10] on top of their backbone to obtain multi-scale features, significantly improving their performance. Overall, without using additional modules (e.g., Atrous Spatial Pyramid Pooling [38], PPM) or their modified versions, our method with moderate model size achieves the highest segmentation accuracy among all real-time methods.
Table 11: Performance on CamVid. We use the whole picture as input to compute mIoU.
@@ 4.3.1 Accuracy Comparison
Table. 10 shows the comparison of accuracy and model complexity with accuracy-oriented methods, most of which use a very deep classification network (e.g., ResNet-101) or a well-designed large-scale network (e.g., HRNet) as the backbone. Our method shows a very competitive result with these state-of-the-art methods with much smaller model complexity. Compared to the most accurate method HRNet-V2, our network reduces the accuracy by less than 3.0% at 19.7% of the model size, 29.2% of the FLOPs and 3.7$\times$ FPS. Furthermore, our network reaches a higher segmentation accuracy with a much smaller model complexity than most methods, including the well-known DeepLab, RefineNet, PSPNet, etc. For example, our model outperforms PSPNet by 0.3% mIoU with 14.6% FLOPs, 19.8% Params and 6.2$\times$ FPS of it. The above results show that our network has a light-weight and effective decoder, and demonstrates this multi-scale parallel inference architecture is one way to achieve efficient semantic segmentation.
4.3.2 Trade-off between Speed and Accuracy
Fig. 1 reflects the trade-off between FPS and accuracy, from which we have the following observations: 1) Our MFARANet and its pruned versions using different input size reach the highest accuracy among all real-time methods. 2) Only STDC2-Seg75 could reach competitive accuracy with our model (only 0.3% lower than us) on the test set at a similar inference speed, while lower than us by 0.9% on the valuation set (Table. 9). Additionally, our pruned-model has the same accuracy as STDC2-Seg75 on the test set, and reaches 0.7% higher accuracy on the valuation set with a speed increase of 12.1 FPS. 3) For methods having a faster inference speed, our MFARANet-pruned at 64.4 FPS achieves 1.2%-9.5% accuracy improvement on the test set. Thus, we can draw the conclusion that our framework exhibits a better trade-off than the compared methods.
4.4 Evaluation on Other Datasets
4.4.1 Performance on CamVid
Table 12: Performance on PASCAL-Context. All results are reported using multi-scale testing. We also report the performance of our model with different backbones.
@@ 4.4.1 Performance on CamVid
CamVid is another dataset of road scene from a car perspective, which has much fewer training and testing samples than Cityscapes. We report the results on it in Table. 11 to prove the generality of our approach, from which we can draw similar conclusions as on the Cityscapes dataset. Compared to speed-oriented methods, our method shows a large increase in accuracy. For example, compared to ENet and CGNet, the accuracy of our method improves by 21.6% and 7.3%, respectively. Among methods aiming at good trade-off, we achieve the highest accuracy with a moderate model size.
4.4.2 Performance on PASCAL-Context
@@
PASCAL-Context is a very challenging scene understanding dataset. Quantitative results are shown in Table. 12, and all comparisons are picked from accuracy-oriented methods. Using ResNet-18 as the encoder, our method can outperform half of the accuracy-oriented methods in the table with very few parameters. By employing ResNet-50, the MFARANet can exceed the performance of all methods except EncNet using ResNet-101, including RefineNet adopting the very deep classification network ResNet-152 as the backbone. In particular, when using the same ResNet-50 as the encoder, our method improves the accuracy by 0.4% and 4% compared to EncNet and ShelfNet50, while having a much smaller model size (25.7M vs. 35.5M and 38.7M). The above results demonstrate the efficiency and effectiveness of our method.
5 Discussion
Our work conducts an in-depth and detailed study of the parallel inference architecture to make it suitable for real-time segmentation tasks. In this section, combined with the experimental results, we summarize the overall performance of MFARANet, and mainly discuss the motivation and substantiation of the three core components proposed for building our parallel inference network. As for the analysis of Multi-scale Joint Supervision and the proof that ASFM can favor objects of multiple scales, we provide them in Section. 4.2.6 and Section. 4.2.7 respectively.
The results in Table. 9 and Table. 11 show that our network achieves the highest accuracy with a moderate model complexity among all real-time methods on the Citysapes and CamVid datasets. The trade-off comparison in Fig. 1 further prove that our model achieves the best balance between performance and model complexity. Besides, we achieve better or competitive performance with much smaller model complexity compared to accuracy-oriented methods in Table. 10 and Table. 12. The above results demonstrate the efficiency of the proposed multi-scale parallel inference network, and show its potential in combination with more advanced backbones for accuracy-oriented methods.
The MFAM, RAM, and ASFM are specially designed for parallel inference segmentation network. Firstly, Table. 2 proves that the MFAM is more suitable for segmentation than FPN because of aggregating multi-level features to each scale to provid hierarchical clues for inference. Secondly, by comparing RAM with bilinear interpolation and Straightforward Alignment, Table. 3 demonstrates that spatial misalignment is one of the key issues for pixel-level scores fusion and can be efficiently alleviated by our RAM. Finally, for ASFM, Table. 4 and Table. 7 show the effectiveness of adaptive fusion and the necessity of the prediction at each scale. Furthermore, the third row of Table. 5 proves that this adaptive score-level fusion is more effective and efficient than the feature-level fusion and single-scale prediction used by most segmentation networks [25, 19, 42].
Our network has some limitations that can be improved. The pixel-level attention mechanism in ASFM helps improve the performance, but it also increases the computational budget of the network. Although the pruning method shown in Fig. 9 and discussed in Section. 4.2.8 largely reduces model complexity, it is highly specific for our parallel inference architecture and its generality need to be further studied. Therefore, designing a fast attention mechanism has the potential to further improve the speed and generality of our networks. Another future work is to build Transformer branches upon the MFAM to explore the combination of this parallel inference architecture with the self-attention mechanism [88].
6 Conclusion